OpenStudy (anonymous):

The number of boys in ninth grade is 12 less than twice the number of girls. There are 51 students in ninth grade. How many are girls?

OpenStudy (anonymous):

X=girls 2x-12 = boys 2x-12+x= 51 3x-12=51 3x-12+12=51+12 3x=63 3x/3=63/3 X =21 21 girls 30 boys